OpenStudy (anonymous):

A triangle has the side lengths 5, 7, and 11. Which term best describes the triangle?

jimthompson5910 (jim_thompson5910):

Notice how 11^2 = 121 is larger than 5^2+7^2 = 25+49=74 So because a^2 + b^2 < c^2, this tells us that we have an obtuse triangle.

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

sure thing